# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ Appendix B Dual representation
In this section, we present the facts concerning the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}(\mathbf{C}^{N})^{*}$. Unless otherwise stated, their proofs mirror the ones for $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$ and will be omitted.
@@
(B.1) $\Omega^{*}\colon\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*},\qquad x\phi_{i}\mapsto\sum_{j}\phi_{j}E_{ji}.$
@@
$(\Omega^{*}(\phi_{1}),\ldots,\Omega^{*}(\phi_{N}))^{T}=(E^{T}+N\hbar\cdot\mathrm{Id})\cdot(\phi_{1},\ldots,\phi_{N})^{T}.$
In particular, the analog of the Cayley-Hamilton theorem of Proposition˜3.7 reads as
@@
which follows directly from [Mol07, Theorem 7.2.1]. Similarly to Proposition˜1.3, one can show that
(B.2) $A(u)^{-1}\phi A(u)=\frac{u-\Omega^{*}}{u-\Omega^{*}+\hbar}\phi.$
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes((\mathbf{C}^{N})^{*})^{\otimes k}\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\ldots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}.$
@@
(B.3) $\begin{split}\Omega^{*}_{i}:=1\otimes\ldots\otimes\Omega^{*}\otimes\ldots\otimes 1\curvearrowright\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\ldots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*},\\ \sigma_{i}(x\otimes\phi_{a_{1}}\otimes\ldots\otimes\phi_{a_{i}}\otimes\phi_{a_{i+1}}\otimes\ldots\otimes\phi_{a_{k}}):=x\otimes\phi_{a_{1}}\otimes\ldots\otimes\phi_{a_{i+1}}\otimes\phi_{a_{i}}\otimes\ldots\otimes\phi_{a_{k}},\end{split}$
Recall the notations for the parabolic restriction of Section˜4. As in (4.10), we can define
(B.4) $\bar{\phi}_{i}=\sum\limits_{1\leq i_{1}<\ldots<i_{s}<i}P(w_{i}-w_{j_{1}})\ldots(w_{i}-w_{j_{r}})\phi_{i_{1}}E_{i_{1}i_{2}}\ldots E_{i_{s}i}\in\mathfrak{n}_{-}\backslash\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*},$
where $s=0,1\ldots$ and $\{j_{1},\ldots,j_{r}\}$ is the complementary subset to $\{i_{1},\ldots,i_{s}\}$ in $\{1\ldots,i-1\}$. By the trick of Example˜3.6 and [Mol02, (3.3)], it is essentially given by the action of the extremal projector
(B.5) $\bar{\phi}_{i}=(w_{i}-w_{1})\ldots(w_{i}-w_{i-1})P\phi_{i}P.$
@@
(B.6) $P\Omega^{*}(\phi_{i})P=(w_{i}+\hbar)P\phi_{i}P\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}/\!\!/N_{-}$
as in ˜4.6.
Recall the positive part $^{N}\tilde{\mathcal{A}}_{\hbar}^{+}$ of the algebra of difference operators $^{N}\tilde{\mathcal{A}}_{\hbar}$ from Section˜4.4. We have an analog of Theorem˜4.11
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}(\mathbf{C}^{N})^{*}/\!\!/N_{-}\xrightarrow{\sim}{}^{N}\tilde{\mathcal{A}}_{\hbar}^{+},\qquad\bar{\phi}_{i}\mapsto\mathbf{u}_{i}.$
Next, recall the Kostant-Whittaker reduction of Section˜5, in particular the Kirillov projector $P_{\psi}$. The following equality can be proved analogously to [Kal25, Proposition 5.20].
(B.7) $u^{N}-\sum_{i=1}^{N}\psi^{N-i}u^{i-1}P_{\psi}E_{Ni}P_{\psi}=P_{\psi}A(u)P_{\psi}.$
The following is an analog of Proposition˜5.5.
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}]/(A(\Omega^{*}-\hbar))\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+},\qquad f\mapsto f(\phi_{N}).$
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes(\mathbf{C}^{N})^{*}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+},\qquad\phi_{i}\mapsto P_{\psi}\phi_{i}P_{\psi}$
is an isomorphism. Therefore, it will be enough to relate the action of $\Omega^{*}$ to that of the Kirillov projector. By pushing the $\mathfrak{gl}_{N}$-terms to the left or to the right and using of Theorem˜5.2 as in [Kal25, Proposition 6.18], we deduce
@@
In particular, the matrix of $\zeta=\Omega^{*}$-$\hbar$ in the generators $\{P_{\psi}\phi_{i}P_{\psi}|i=1\ldots N\}$ has the form
$\begin{pmatrix}0&\psi&\ldots&0&0\\ 0&0&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\\ 0&0&\ldots&0&\psi\\ -\psi^{-N+1}A_{N}&-\psi^{-N+2}A_{N-1}&\ldots&-\psi^{-1}A_{2}&-A_{1}\end{pmatrix}$
Define polynomials $Q_{i}(\zeta)\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\zeta]$ for $i=1\ldots N$ inductively by the rule
$Q_{N}(\zeta)=1,\qquad Q_{i-1}(\zeta)=\psi^{-1}[\psi^{N-i}A_{N-i}+\zeta Q_{i}(\zeta)].$
The transition matrix from $(1,\ldots,\zeta^{N-1})$ to $(\psi_{N}(\zeta),\ldots,\psi_{1}(\zeta))$ is upper triangular with invertible diagonal, hence is invertible itself. However, observe that the action of $\zeta$ on the latter polynomials is given by the matrix above. In particular, since the elements $\{P_{\psi}\phi_{i}P_{\psi}|i=1\ldots N\}$ generate $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+}$ as a left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-module, the elements $\{(\Omega^{*}-\hbar)^{i-1}(\phi_{N})|i=1\ldots N\}$ do as well. However, the latter generate $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}]/(A(\Omega^{*}-\hbar))$. ∎
@@
with values in $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}_{1},\ldots,\Omega^{*}_{k}]$ as in (5.5). Likewise, consider intersection
$I^{*}_{k}:=(A^{(-j)}(\Omega_{j}^{*}-\hbar)|j=1\ldots k)\cap\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}_{1},\ldots,\Omega^{*}_{k}].$
The following is an analog of Proposition˜5.7 and ˜5.8.
@@
    $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}_{1},\ldots,\Omega^{*}_{k}]/(A^{(-j)}(\Omega^{*}_{j}-\hbar)|j=1\ldots k)\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes((\mathbf{C}^{N})^{*})^{\otimes k}/\!\!/_{\psi}\,N_{+},$
@@
    $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}_{1},\ldots,\Omega^{*}_{k}]^{S_{k}}/I^{*}_{k}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+}.$
Recall the definition of the non-positive part $\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}$ of the shifted Yangian from ˜2.13. The following is an analog of Theorem˜5.9.
@@
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+},$
@@
$x_{-}(u)$ $\mapsto\frac{1}{u-\Omega^{*}_{1}}(\phi_{N})\in(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+})[[u^{-1}]].$
Finally, recall the Miura transform of ˜6.4:
(B.8) $\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}(\mathbf{C}^{N})^{*}/\!\!/N_{-}\cong{}^{N}\tilde{\mathcal{A}}_{\hbar}^{-}$
Using (B.6), the following theorem can be proved analogously to Theorem˜6.7.
@@
The map (B.8) coincides with the GKLO map from Theorem˜2.16.
The following result is used in Section˜7.2.
@@
Under the embedding $\Lambda^{k}(\mathbf{C}^{N})^{*}\hookrightarrow((\mathbf{C}^{N})^{*})^{\otimes k}$, we have
$\phi_{N-k+1}\wedge\ldots\wedge\phi_{N}P_{\psi}=\frac{\psi^{-\frac{k(k-1)}{2}}}{k!}\left[\prod_{i<j}(\Omega_{i}^{*}-\Omega_{j}^{*}+\hbar)\right](\phi_{N}^{\otimes k})P_{\psi},$
where both sides are in $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\Lambda^{k}(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+}$.
@@
Since $\phi_{N-k+1}\wedge\ldots\wedge\phi_{N}$ is a highest-weight vector, we have
$\phi_{N-k+1}\wedge\ldots\wedge\phi_{N}P_{\psi}=P_{\psi}\phi_{N-k+1}\wedge\ldots\wedge\phi_{N}P_{\psi}.$
@@
$P_{\psi}\phi_{N-k+1}\wedge\ldots\wedge\phi_{N}P_{\psi}=e^{-}_{k}\cdot P_{\psi}\phi_{N-k+1}\otimes\ldots\otimes\phi_{N}P_{\psi}=\psi^{-\frac{k(k-1)}{2}}(e^{-}_{k})^{-\hbar}[(\Omega^{*}_{1})^{k-1}\ldots\Omega^{*}_{k-1}](\phi_{N}^{\otimes k}).$
where $e^{-}_{k}$ is the antisymmetrizer (2.6) acting on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes((\mathbf{C}^{N})^{*})^{\otimes k}$, hence on its Kostant-Whittaker reduction. The case $k=1$ is obvious, and for the step, it follows from the proof of Proposition˜B.1 that
@@
$e^{-}_{k}P_{\psi}\phi_{N-k+1}\otimes\ldots\otimes\phi_{N}P_{\psi}=\psi^{-\frac{k(k-1)}{2}}e^{-}_{k}P_{\psi}(\Omega^{*}_{1})^{k-1}\ldots\Omega^{*}_{k-1}(\phi_{N}^{\otimes k})P_{\psi}$
By $\mathfrak{H}^{-\hbar}_{k}$-equivariance of Proposition˜B.1 and Proposition˜2.7, the right-hand side is equal to the action of the element
$\frac{\psi^{-\frac{k(k-1)}{2}}}{k!}\prod_{i<j}\frac{Y_{i}-Y_{j}+\hbar}{Y_{i}-Y_{j}}\cdot\mathrm{Sym}_{k}^{-}(Y_{1}^{k-1}\ldots Y_{k-1}^{1}Y_{k}^{0})=\frac{\psi^{-\frac{k(k-1)}{2}}}{k!}\prod_{i<j}(Y_{i}-Y_{j}+\hbar)$
@@
$e^{-}_{k}\cdot(\Omega^{*}_{1})^{k-1}\ldots\Omega^{*}_{k-1}(\phi_{N}^{\otimes k})P_{\psi}=\frac{\psi^{-\frac{k(k-1)}{2}}}{k!}\left[\prod_{i<j}(\Omega_{i}^{*}-\Omega_{j}^{*}+\hbar)\right](\phi_{N}^{\otimes k})P_{\psi},$
@@
Finally, consider the topological setting of Section˜7.1. Let
$\mathrm{Gr}^{\omega_{1}}=\overline{\mathrm{Gr}^{\omega_{1}}}=\{zL_{0}\subset L_{-1}\subset L_{0}|\dim L_{0}/L_{-1}=1\}.$
There is a factorization similar to (7.10):
(B.9) ${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varpi}_{N}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes\mathbf{C}^{N}/\!\!/_{\psi^{*}}\,N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}},\underline{\mathbf{C}}_{\mathrm{Gr}^{\omega_{1}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})}$
@@
$A^{(i-1)}(u)=\sum_{j=0}^{N}u^{N-j}A_{j}^{(i-1)},\qquad A_{j}^{(i-1)}=(-1)^{i}c_{j}^{\mathrm{GL}_{N}}(\mathcal{V}_{i}).$
@@
The following is an analog of Proposition˜7.13. Recall the isomorphism of Proposition˜B.1. Also, since the loop rotation is trivial on $\mathcal{Q}$, we have
(B.10) $c_{1}^{G}(\mathcal{Q})=c_{1}^{\mathrm{GL}_{N}}(\mathcal{Q})$

### v4

@@ Appendix B Dual representation
In this appendix, we present the facts concerning the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}$. Unless otherwise stated, their proofs mirror the ones for $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$ and will be omitted.
@@
$\Omega^{*}\colon\ \mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*},\qquad x\phi_{i}\mapsto\sum_{j}\phi_{j}E_{ji}.$ (B.1)
@@
$(\Omega^{*}(\phi_{1}),\dots,\Omega^{*}(\phi_{N}))^{\mathsf{T}}=\bigl(E^{\mathsf{T}}+N\hbar\cdot\mathrm{Id}\bigr)\cdot(\phi_{1},\dots,\phi_{N})^{\mathsf{T}}.$
In particular, the analog of the Cayley–Hamilton theorem of Proposition 4.7 reads as
@@
which follows directly from [38, Theorem 7.2.1]. Similarly to Proposition 2.3, one can show that
$A(u)^{-1}\phi A(u)=\frac{u-\Omega^{*}}{u-\Omega^{*}+\hbar}\phi.$ (B.2)
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)^{\otimes k}\cong\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\dots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}.$
@@
$\Omega^{*}_{i}:=1\otimes\dots\otimes\Omega^{*}\otimes\dots\otimes 1\curvearrowright\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\dots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*},$
$\sigma_{i}(x\otimes\phi_{a_{1}}\otimes\dots\otimes\phi_{a_{i}}\otimes\phi_{a_{i+1}}\otimes\dots\otimes\phi_{a_{k}}):=x\otimes\phi_{a_{1}}\otimes\dots\otimes\phi_{a_{i+1}}\otimes\phi_{a_{i}}\otimes\dots\otimes\phi_{a_{k}}.$
Recall the notations for the parabolic restriction of Section 5. As in (5.9), we can define
$\bar{\phi}_{i}=\sum\limits_{1\leq i_{1}<\dots<i_{s}<i}P(w_{i}-w_{j_{1}})\cdots(w_{i}-w_{j_{r}})\phi_{i_{1}}E_{i_{1}i_{2}}\cdots E_{i_{s}i}\in\mathfrak{n}_{-}\backslash\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*},$ (B.3)
where $s=0,1,\dots$ and $\{j_{1},\dots,j_{r}\}$ is the complementary subset to $\{i_{1},\dots,i_{s}\}$ in $\{1\dots,i-1\}$. By the trick of Example 4.6 and [37, formula (3.3)], it is essentially given by the action of the extremal projector
$\bar{\phi}_{i}=(w_{i}-w_{1})\cdots(w_{i}-w_{i-1})P\phi_{i}P.$ (B.4)
@@
$P\Omega^{*}(\phi_{i})P=(w_{i}+\hbar)P\phi_{i}P\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/N_{-}$ (B.5)
as in Lemma 5.6.
Recall the positive part $^{N}\!\tilde{\mathcal{A}}_{\hbar}^{+}$ of the algebra of difference operators $^{N}\!\tilde{\mathcal{A}}_{\hbar}$ from Section 5.4. We have an analog of Theorem 5.11
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/N_{-}\xrightarrow{\sim}{}^{N}\!\tilde{\mathcal{A}}_{\hbar}^{+},\qquad\bar{\phi}_{i}\mapsto\mathbf{u}_{i}.$
Next, recall the Kostant–Whittaker reduction of Section 6, in particular the Kirillov projector $P_{\psi}$. The following equality can be proved analogously to [27, Proposition 5.20].
$u^{N}-\sum_{i=1}^{N}\psi^{N-i}u^{i-1}P_{\psi}E_{Ni}P_{\psi}=P_{\psi}A(u)P_{\psi}.$ (B.6)
The following is an analog of Proposition 6.5.
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}]/(A(\Omega^{*}-\hbar))\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+},\qquad f\mapsto f(\phi_{N}).$
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+},\qquad\phi_{i}\mapsto P_{\psi}\phi_{i}P_{\psi}$
is an isomorphism. Therefore, it will be enough to relate the action of $\Omega^{*}$ to that of the Kirillov projector. By pushing the $\mathfrak{gl}_{N}$-terms to the left or to the right and using of (6.2) as in [27, Proposition 6.18], we deduce
@@
In particular, the matrix of $\zeta=\Omega^{*}$-$\hbar$ in the generators $\{P_{\psi}\phi_{i}P_{\psi}\mid i=1,\dots,N\}$ has the form
$\begin{pmatrix}0&\psi&\dots&0&0\\ 0&0&\dots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\\ 0&0&\dots&0&\psi\\ -\psi^{-N+1}A_{N}&-\psi^{-N+2}A_{N-1}&\dots&-\psi^{-1}A_{2}&-A_{1}\end{pmatrix}$
Define polynomials $Q_{i}(\zeta)\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\zeta]$ for $i=1,\dots,N$ inductively by the rule
$Q_{N}(\zeta)=1,\qquad Q_{i-1}(\zeta)=\psi^{-1}\bigl[\psi^{N-i}A_{N-i}+\zeta Q_{i}(\zeta)\bigr].$
The transition matrix from $\bigl(1,\dots,\zeta^{N-1}\bigr)$ to $(\psi_{N}(\zeta),\dots,\psi_{1}(\zeta))$ is upper triangular with invertible diagonal, hence is invertible itself. However, observe that the action of $\zeta$ on the latter polynomials is given by the matrix above. In particular, since the elements $\{P_{\psi}\phi_{i}P_{\psi}\mid i=1,\dots,N\}$ generate $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+}$ as a left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-module, the elements $\bigl\{(\Omega^{*}-\hbar)^{i-1}(\phi_{N})\mid$ $i=1,\dots,N\bigr\}$ do as well. However, the latter generate $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}]/(A(\Omega^{*}-\hbar))$. ∎
@@
with values in $\mathbf{Z}_{\mathfrak{gl}_{N}}\bigl[\Omega^{*}_{1},\dots,\Omega^{*}_{k}\bigr]$ as in (6.5). Likewise, consider intersection
$I^{*}_{k}:=\bigl(A^{(-j)}(\Omega_{j}^{*}-\hbar)\mid j=1\dots k\bigr)\cap\mathbf{Z}_{\mathfrak{gl}_{N}}\bigl[\Omega^{*}_{1},\dots,\Omega^{*}_{k}\bigr].$
The following is an analog of Proposition 6.7 and Corollary 6.8.
@@
    $\mathbf{Z}_{\mathfrak{gl}_{N}}\bigl[\Omega^{*}_{1},\dots,\Omega^{*}_{k}\bigr]/\bigl(A^{(-j)}(\Omega^{*}_{j}-\hbar)\mid j=1,\dots,k\bigr)\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+},$
@@
    $\mathbf{Z}_{\mathfrak{gl}_{N}}\bigl[\Omega^{*}_{1},\dots,\Omega^{*}_{k}\bigr]^{S_{k}}/I^{*}_{k}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+}.$
Recall the definition of the non-positive part $\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}$ of the shifted Yangian from Definition 3.13. The following is an analog of Theorem 6.9.
@@
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+},$
@@
$x_{-}(u)$ $\mapsto\frac{1}{u-\Omega^{*}_{1}}(\phi_{N})\in\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+}\bigr)\bigl[\hskip-0.56905pt\bigl[u^{-1}\bigr]\hskip-0.56905pt\bigr].$
Finally, recall the Miura transform of Definition 7.4:
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/N_{-}\cong{}^{N}\!\tilde{\mathcal{A}}_{\hbar}^{-}.$ (B.7)
Using (B.5), the following theorem can be proved analogously to Theorem 7.7.
@@
The map (B.7) coincides with the GKLO map from Theorem 3.16.
The following result is used in Section 8.2.
@@
Under the embedding $\Lambda^{k}\bigl(\mathbf{C}^{N}\bigr)^{*}\hookrightarrow\bigl(\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)^{\otimes k}$, we have
$\phi_{N-k+1}\wedge\dots\wedge\phi_{N}P_{\psi}=\frac{\psi^{-\frac{k(k-1)}{2}}}{k!}\biggl[\prod_{i<j}\big(\Omega_{i}^{*}-\Omega_{j}^{*}+\hbar\big)\biggr]\bigl(\phi_{N}^{\otimes k}\bigr)P_{\psi},$
where both sides are in $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\Lambda^{k}\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+}$.
@@
Since $\phi_{N-k+1}\wedge\dots\wedge\phi_{N}$ is a highest-weight vector, we have
$\phi_{N-k+1}\wedge\dots\wedge\phi_{N}P_{\psi}=P_{\psi}\phi_{N-k+1}\wedge\dots\wedge\phi_{N}P_{\psi}.$
@@
$P_{\psi}\phi_{N-k+1}\wedge\dots\wedge\phi_{N}P_{\psi}$ ${}={\rm e}^{-}_{k}\cdot P_{\psi}\phi_{N-k+1}\otimes\dots\otimes\phi_{N}P_{\psi}$
${}=\psi^{-\frac{k(k-1)}{2}}({\rm e}^{-}_{k})^{-\hbar}\bigl[(\Omega^{*}_{1})^{k-1}\cdots\Omega^{*}_{k-1}\bigr]\bigl(\phi_{N}^{\otimes k}\bigr),$
where ${\rm e}^{-}_{k}$ is the antisymmetrizer (3.6) acting on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)^{\otimes k}$, hence on its Kostant–Whittaker reduction. The case $k=1$ is obvious, and for the step, it follows from the proof of Proposition B.1 that
@@
${\rm e}^{-}_{k}P_{\psi}\phi_{N-k+1}\otimes\dots\otimes\phi_{N}P_{\psi}=\psi^{-\frac{k(k-1)}{2}}{\rm e}^{-}_{k}P_{\psi}(\Omega^{*}_{1})^{k-1}\cdots\Omega^{*}_{k-1}\bigl(\phi_{N}^{\otimes k}\bigr)P_{\psi}.$
By $\mathfrak{H}^{-\hbar}_{k}$-equivariance of Propositions B.1 and 3.7, the right-hand side is equal to the action of the element
$\frac{\psi^{-\frac{k(k-1)}{2}}}{k!}\prod_{i<j}\frac{Y_{i}-Y_{j}+\hbar}{Y_{i}-Y_{j}}\cdot\mathrm{Sym}_{k}^{-}\bigl(Y_{1}^{k-1}\cdots Y_{k-1}^{1}Y_{k}^{0}\bigr)=\frac{\psi^{-\frac{k(k-1)}{2}}}{k!}\prod_{i<j}(Y_{i}-Y_{j}+\hbar)$
@@
${\rm e}^{-}_{k}\cdot(\Omega^{*}_{1})^{k-1}\cdots\Omega^{*}_{k-1}\bigl(\phi_{N}^{\otimes k}\bigr)P_{\psi}=\frac{\psi^{-\frac{k(k-1)}{2}}}{k!}\biggl[\prod_{i<j}\big(\Omega_{i}^{*}-\Omega_{j}^{*}+\hbar\big)\biggr]\bigl(\phi_{N}^{\otimes k}\bigr)P_{\psi},$
@@
Finally, consider the topological setting of Section 8.1. Let
$\mathrm{Gr}^{\omega_{1}}=\overline{\mathrm{Gr}^{\omega_{1}}}=\{zL_{0}\subset L_{-1}\subset L_{0}\mid\dim L_{0}/L_{-1}=1\}.$
There is a factorization similar to (8.10),
[…]
@@
$A^{(i-1)}(u)=\sum_{j=0}^{N}u^{N-j}A_{j}^{(i-1)},\qquad A_{j}^{(i-1)}=(-1)^{i}c_{j}^{\mathrm{GL}_{N}}(\mathcal{V}_{i})$
@@
The following is an analog of Proposition 8.14. Recall the isomorphism of Proposition B.1. Also, since the loop rotation is trivial on $\mathcal{Q}$, we have
$c_{1}^{G}(\mathcal{Q})=c_{1}^{\mathrm{GL}_{N}}(\mathcal{Q}).$ (B.9)